The president of a company was interested in determining whether there is a correlation between sales made by different sales teams and hours spent on employee training. These figures are shown.

Sales
(in thousands)

Training
Hours

14 6
33 12
20 11
44 15
8 5
  1. Compute the correlation coefficient for the data. (Round your answer to 4 decimal places, the tolerance is +/-0.0001.)
    The correlation coefficient is

    Entry field with correct answer



    What is your interpretation of this value? (Do not round your intermediate computations to answer this question.)
    There is a strong positive linear association between sales and training hours.
  2. Using the data, what would you expect sales to be if training was increased to eighteen hours? Use the linear regression model. (Round your answer to 2 decimal places, the tolerance is +/-0.01.)
    Sales = (in thousands)
Answer #1

The table given below ,

Training Hours (X) Sales(Y) X^2 Y^2 XY
6 14 36 196 84
12 33 144 1089 396
11 20 121 400 220
15 44 225 1936 660
5 8 25 64 40
\sum X=49 \sum Y=119 \sum X^2=551 \sum Y^2=3685 \sum XY=1400

From table ,

\bar{X}=\frac{\sum X}{n}=\frac{49}{5}=9.8

\bar{Y}=\frac{\sum Y}{n}=\frac{119}{5}=23.8

\sigma^2_{X}=\frac{\sum X^2}{n}-\bar{X}^2=\frac{551}{5}-9.8^2=14.16

\sigma^2_{Y}=\frac{\sum Y^2}{n}-\bar{Y}^2=\frac{3685}{5}-23.8^2=170.56

Cov(X,Y)=\frac{\sum XY}{n}-\bar{X}*\bar{Y}=\frac{1400}{5}-9.8*23.8=46.76

Therefore ,

a) The correlation coefficient for the data is ,

r=\frac{Cov(X,Y)}{\sqrt{\sigma_{X}^2*\sigma_{Y}^2}}=\frac{46.76}{\sqrt{14.16*170.56}}=0.9515

Interpretation : Here , the value of the correlation coefficient is positive and close to 1.

Therefore , there is strong positive linear association between training hours and sales.

b) The regression coefficient is ,

b=\frac{cov(X,Y)}{\sigma_{X}^2}=\frac{46.76}{14.16}=3.3023

The intercept of the regression equation is ,

a=\bar{Y}-b*\bar{X}=23.8-3.3023*9.8=-8.5621

Therefore , the linear regression equation is ,

\hat{y}=a+bx=-8.5621+3.3023x

Now , for x=8 hours

\therefore \hat{y}=-8.5621+3.3023x=-8.5612+3.3023*18=50.88

Therefore , if the training was increased to 18 hours , then the expected sale is 50.88 in thosands.